# Density Ratio Estimation-based Bayesian Optimization with Semi-Supervised Learning

### v4

@@ Appendix A Additional Comparisons of BORE and LFBO
Appendix B Details of DRE-BO-SSL
@@ Appendix C Analysis of DRE-BO-SSL
Using Definition C.1, we claim that a semi-supervised classifier in DRE-BO-SSL can mitigate the over-exploitation problem presented in Section 1.1, because it can expand a decision set $\mathcal{C}_{1}$ for Class 1 by reducing $\gamma$ with unlabeled data. However, we need to verify if a large decision set is derived from the characteristics of non-parametric classifiers, since a semi-supervised classifier we use is converted to the Nadaraya-Watson non-parametric model (Nadaraya, 1964; Watson, 1964) without unlabeled data. As shown in Figure 12, there is no strong relationship between the performance of semi-supervised classifiers without unlabeled data, i.e., the Nadaraya-Watson estimator, and one of semi-supervised classifiers with unlabeled data. It implies that configuration selection for a semi-supervised classifier is dependent on a class of objective function, and the presence of unlabeled data is likely to be effective for alleviating the over-exploitation problem.
@@
Appendix D Experiment Details
@@
To reduce the computational complexity of DRE-BO-SSL for a scenario with a fixed-size pool, we randomly sample 2,000 unlabeled points from the predefined pool for all experiments excluding synthetic benchmark functions. More thorough analysis can be found in Section G.
@@
simple $\textrm{simple regret}(f(\mathbf{x}_{1}),\ldots,f(\mathbf{x}_{t}),f(\mathbf{x}^{*}))=\min_{i=1}^{t}f(\mathbf{x}_{i})-f(\mathbf{x}^{*}),$ (17)
@@
D.1 Details of Synthetic Benchmarks
@@
D.2 Details of Tabular Benchmarks
@@
The search space of Tabular Benchmarks (Klein & Hutter, 2019) is described in Table 1. To handle categorical and discrete variables, we treat them as integer variables by following the previous literature (Garrido-Merchán & Hernández-Lobato, 2020).
D.3 Details of NATS-Bench
@@
D.4 Details of Minimum Multi-Digit MNIST Search
@@
Since multi-digit MNIST, which is composed of images of size $(28,84)$ and shown in Figure 10, is high-dimensional, some of the methods used in this work, e.g., methods with random forests, gradient boosting, and XGBoost, struggle to process such data. Therefore, we embed an original image to a lower-dimensional vector using an auxiliary convolutional neural network. The convolutional neural network is trained to classify a three-digit image to one of labels from “000” to “999,” with the following architecture:
  1. First layer: convolutional, input channel $1$, output channel $8$, kernel size $3\times 3$, padding 1, ReLU, max-pooling $2\times 2$;
  2. Second layer: convolutional, input channel $8$, output channel $16$, kernel size $3\times 3$, padding 1, ReLU, max-pooling $2\times 2$;
  3. Third layer: convolutional, input channel $16$, output channel $32$, kernel size $3\times 3$, padding 1, ReLU, max-pooling $2\times 2$;
  4. Fourth layer: fully-connected, input dimensionality $960$, output dimensionality $128$, ReLU;
  5. Fifth layer: fully-connected, input dimensionality $128$, output dimensionality $64$, ReLU;
  6. Sixth layer: fully-connected, input dimensionality $64$, output dimensionality $1000$, Softmax.
The Adam optimizer (Kingma & Ba, 2015) with learning rate $1\times 10^{-3}$ is used to train the network for 100 epochs. To train and test the model fairly, we create a training dataset of 440,000 three-digit images, a validation dataset of 40,000 three-digit images, and a test dataset of 80,000 three-digit images using a training dataset of 55,000 single-digit images, a validation dataset of 5,000 single-digit images, and a test dataset of 10,000 single-digit images in the original MNIST dataset (LeCun et al., 1998). For example, supposing that a test dataset has 1,000 single-digit images per class—it is not true for the MNIST dataset, but it is assumed for explanation—and we would like to generate a three-digit image “753,” $10^{9}$ combinations for “753” can be created. We therefore randomly sample a fixed number of three-digit images from a vast number of possible combinations. In addition, an early stopping technique is utilized by comparing the current validation loss to the average of validation losses for the recent five epochs. Eventually, our network achieves $99.6\%$ in the training dataset, $97.0\%$ in the validation dataset, and $96.9\%$ in the test dataset.
@@ 5. Fifth layer: fully-connected, input dimensionality $128$, output dimensionality $64$, ReLU;
Appendix E Discussion on a Free Parameter in Label Propagation and Label Spreading
@@
Appendix F Discussion on Unlabeled Point Sampling
@@
Appendix G Discussion on Pool Sampling
@@
Appendix H Discussion on a Threshold Ratio
@@
Figure 16 demonstrates the effects of a threshold ratio $\zeta$, where we use $\beta=0.5$. It follows the setting described in Section E. As presented in Figure 16, a smaller $\zeta$, for example, $\zeta=0.01$ or $\zeta=0.1$ tends to show worse performance than a larger $\zeta$, which implies that the over-exploitation is not due to conservative $y^{\dagger}$. Interestingly, the results with $\zeta=0.8$ also generally under-perform. We presume that it is basically due to over-exploration.
Appendix I Limitations

### v5

@@ Appendix A Additional Comparisons of BORE and LFBO
Appendix B Additional Related Work
Although many recent Bayesian optimization methods using regression models (Eriksson et al., 2019; Balandat et al., 2020; Cowen-Rivers et al., 2022; Ament et al., 2023) show strong performance on various benchmark functions, we do not compare our methods to them, as such comparisons are beyond the scope of this work. The goal of this work is to demonstrate the effectiveness of DRE-based Bayesian optimization with semi-supervised learning in the settings of DRE-based Bayesian optimization. Picheny et al. (2019) propose a novel Bayesian optimization framework that relies on variable ordering in a latent space to handle ill-conditioned or discontinuous objectives. Regarding the choice of the truncated multivariate normal distribution, it may be related to the boundary issue of Bayesian optimization; see the Ph.D. thesis of Swersky (2017). Such a similar issue is discussed in the existing work by Oh et al. (2018); it proposes a Bayesian optimization approach with cylindrical kernels concentrating on a region proximal to the center of the search space. Moreover, the prior work by Hvarfner et al. (2024) also reports the similar consequence in Bayesian optimization. From the perspective of DRE-based Bayesian optimization, Gaussian process classification (Rasmussen & Williams, 2006) might be applicable for defining binary classifiers; it is left for future work.
Appendix C Details of DRE-BO-SSL
@@ Appendix D Analysis of DRE-BO-SSL
Using Definition D.1, we claim that a semi-supervised classifier in DRE-BO-SSL can mitigate the over-exploitation problem presented in Section 1.1, because it can expand a decision set $\mathcal{C}_{1}$ for Class 1 by reducing $\gamma$ with unlabeled data. However, we need to verify if a large decision set is derived from the characteristics of non-parametric classifiers, since a semi-supervised classifier we use is converted to the Nadaraya-Watson non-parametric model (Nadaraya, 1964; Watson, 1964) without unlabeled data. As shown in Figure 12, there is no strong relationship between the performance of semi-supervised classifiers without unlabeled data, i.e., the Nadaraya-Watson estimator, and one of semi-supervised classifiers with unlabeled data. It implies that configuration selection for a semi-supervised classifier is dependent on a class of objective function, and the presence of unlabeled data is likely to be effective for alleviating the over-exploitation problem.
@@
Appendix E Experimental Details
@@
To reduce the computational complexity of DRE-BO-SSL for a scenario with a fixed-size pool, we randomly sample 2,000 unlabeled points from the predefined pool for all experiments excluding synthetic benchmark functions. More thorough analysis can be found in Section H.
@@
$\textrm{simple regret}(f(\mathbf{x}_{1}),\ldots,f(\mathbf{x}_{t}),f(\mathbf{x}^{*}))=\min_{i=1}^{t}f(\mathbf{x}_{i})-f(\mathbf{x}^{*}),$ (17)
@@
E.1 Details of Synthetic Benchmarks
@@
E.2 Details of Tabular Benchmarks
@@
The search space of Tabular Benchmarks (Klein & Hutter, 2019) is described in Table 1. To handle categorical and discrete variables, we treat them as integer variables by following the previous literature by Garrido-Merchán & Hernández-Lobato (2020). More precisely, the value of each integer variable corresponds to the index of the original categorical or discrete variable. When evaluating the variables, the integer variables are transformed into the original variables.
E.3 Details of NATS-Bench
@@
E.4 Details of Minimum Multi-Digit MNIST Search
@@
Since multi-digit MNIST, which is composed of images of size (28,84)(28,84)( 28 , 84 ) and shown in Figure 10, is high-dimensional, some of the methods used in this work, e.g., methods with random forests, gradient boosting, and XGBoost, struggle to process such data. Therefore, we embed an original image to a lower-dimensional vector using an auxiliary convolutional neural network. The convolutional neural network is trained to classify a three-digit image to one of labels from “000” to “999,” with the following architecture:
  1. First layer: convolutional, input channel 111, output channel 888, kernel size $\times 3$ 3, padding 1, ReLU, max-pooling $\times 2$ 2;
  2. Second layer: convolutional, input channel 888, output channel 161616, kernel size $\times 3$ 3, padding 1, ReLU, max-pooling $\times 2$ 2;
  3. Third layer: convolutional, input channel 161616, output channel 323232, kernel size $\times 3$ 3, padding 1, ReLU, max-pooling $\times 2$ 2;
  4. Fourth layer: fully-connected, input dimensionality 960960960, output dimensionality 128128128, ReLU;
  5. Fifth layer: fully-connected, input dimensionality 128128128, output dimensionality 646464, ReLU;
  6. Sixth layer: fully-connected, input dimensionality 646464, output dimensionality 100010001000, Softmax.
[…]
@@ 5. Fifth layer: fully-connected, input dimensionality 128128128, output dimensionality 646464, ReLU;
Appendix F Discussion on a Free Parameter in Label Propagation and Label Spreading
@@
Appendix G Discussion on Unlabeled Point Sampling
@@
Appendix H Discussion on Pool Sampling
@@
Appendix I Discussion on a Threshold Ratio
@@
Figure 16 demonstrates the effects of a threshold ratio $\zeta$, where we use $\beta=0.5$. It follows the setting described in Section F. As presented in Figure 16, a smaller $\zeta$, for example, $\zeta=0.01$ or $\zeta=0.1$ tends to show worse performance than a larger $\zeta$, which implies that the over-exploitation is not due to conservative $y^{\dagger}$. Interestingly, the results with $\zeta=0.8$ also generally under-perform. We presume that it is basically due to over-exploration.
Appendix J Limitations